OpenStudy (anonymous):

Find the series interval of convergence and within this interval the sum of the series as a function of x

OpenStudy (anonymous):

\[\sum_{0}^{\infty} (x+1)^{2n}/9^{n}\]

OpenStudy (anonymous):

Can I integrate? hehe

OpenStudy (anonymous):

this is continuous function right?

OpenStudy (anonymous):

but i think it would be a waste to integrate

OpenStudy (anonymous):

how to do write the "over" division line without it including summation

OpenStudy (anonymous):

Hmm \frac{}{} <<use this

OpenStudy (anonymous):

when you said integrate.. what did you mean?

OpenStudy (anonymous):

Hmm that was kinds foolish I guess umm I was talking about integrating the function from o to infinity but i don't think it's integration is possible

OpenStudy (anonymous):

is that something to do with taylor series? because i have not learned that yet.

OpenStudy (anonymous):

taylor series hmm it's about differentiating the function about zero hmm i have a good video link for it lemme fetch that for you

OpenStudy (anonymous):

http://www.khanacademy.org/video/taylor-series-at-0--maclaurin--for-e-to-the-x?playlist=Calculus try this :-D i learned about taylor series from khan academy only but somehow i don't remember much of it, poor memory :/

OpenStudy (anonymous):

you don't have to taylor series. It's the root test first to get the interval of convergence.

OpenStudy (anonymous):

\[\sum_{0}^{\infty} \frac{}{}\ \left( (x+1)^2\over9 \right)^n\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

|r| <1 could i just solve for the intervals of convergence?

OpenStudy (anonymous):

soo.. ∑0∞(x+1)2n/9n = \[\sum_{1}^{\infty}\left| (an+1) \div an \right| = L < 1\]

OpenStudy (anonymous):

-4<x<2 abs convergent because it would be geometric progression, i thought the root test was |an|^1/n

OpenStudy (anonymous):

ok then you test the two endpoints. x=-4 x=2 for absolute convergence

OpenStudy (anonymous):

is that okay? im still unsure of myself with power series

OpenStudy (anonymous):

and to sum the series as a function of x, just use a1/1-r? where r = (1/9)(x-1)^2

OpenStudy (anonymous):

well thanks for you help franklin

OpenStudy (anonymous):

its getting hard to focus.. thanks for your help ishaan

OpenStudy (anonymous):

yea sorry I was trying to find my old notes, forgot how to rewrite the sumation

OpenStudy (anonymous):

sleep is calling

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

stupid even numbered problems :)